OpenStudy (anonymous):

Find any stationary values of thhe following curves and determine whether they are maxima or minima. Sketch the curves: (a) y=x-Inx (b) y=1/2x^2-In2x

OpenStudy (mimi_x3):

Find the first derivative. Then set it to 0 to find the stationary point. To determine if it's a maxima or minima find the second derivative of the function.

OpenStudy (anonymous):

lol what she said^ except when you say maxima and minima, do you mean the maximum / minimum of f(x) or the maximum / minimum of f ' (x)

OpenStudy (anonymous):

Ok, but how do you find the derivative of (a) for instance. I can find it when it's just Inx, but not when there's x-Inx.

OpenStudy (mimi_x3):

\[\large y' = 1-\frac{1}{x} =>\frac{x-1}{x} \]

OpenStudy (mimi_x3):

since derivative of x is 1.

OpenStudy (anonymous):

How did you get to that answer?

OpenStudy (anonymous):

And how would you get the second derivative?

OpenStudy (mimi_x3):

Get to what answer? the derivative of x is 1 and the derivative of lnx is 1/x. Then set it to zero to find the stationary point.

OpenStudy (anonymous):

Well x - lnx, if you want to differentiate, you just do it like any function: d/dx (x - lnx) = d/dx (x) - d/dx(lnx)

OpenStudy (anonymous):

Um, Ok... I understand that now, @Mimi, you said that To determine if it's a maxima or minima find the second derivative of the function. So, how do you get the second derivative?

OpenStudy (anonymous):

Just derive it again: d/dx( 1 - 1/x) = 0 + 1/x^2 = 1/x^2 But i'm pretty sure you don't need the second derivative. I think you can just set the first derivative to 0: 0 = 1 - 1/x x = 1 so plug this back into your original function: y = 1 - ln(1) y = 1 - 0 y = 1 so you're point is (1,1) and it must be a minimum since the range of the function is 1 to infinity.

OpenStudy (mimi_x3):

I think you need to find the second derivative to determine the nature, like if it's a minimum or maximum.

OpenStudy (anonymous):

Ok. Thanks! I understand now. I'm not ompletely sure of deriving still... What's the formula for deriving?

OpenStudy (anonymous):

how would you derive (x + x^2)?

OpenStudy (anonymous):

1 + 2x? Not sure

OpenStudy (mimi_x3):

Use the formula: \[\large \frac{d}{dx} (ax+b)^{n} = n(ax+b)^{n-1} *a\]

OpenStudy (anonymous):

you're right! For derivatives, if you're adding/subtracting them, then you can just derive them individually so for (x - lnx), you find the derivative of (x) which is 1, and you minus the derivative of lnx, which is 1/x y = (x - lnx) d/dx (y) = d/dx (x - lnx) d/dx (y) = d/dx (x) - d/dx (lnx) d/dx (y) = 1 - 1/x

OpenStudy (anonymous):

But, how do you work with derivatives of In? That's what I'm stuck on

OpenStudy (mimi_x3):

The derivative of lnx => 1/x

OpenStudy (mimi_x3):

You just need to remember it.

OpenStudy (anonymous):

derivative of lnx = 1/x I mean, if you want to see the proof for it, I could probably type it up... but it's really long

OpenStudy (anonymous):

No, that's fine... The Inx is fine, it's just, what if it was 2In4x and you had to derive that?

OpenStudy (anonymous):

well then that just becomes 2(1/4x)(4) = 8(1/4x) = 8/4x = 2/x you have to chain rule the 4x, which will pop out a 4.

OpenStudy (anonymous):

Ok.. How about In 1/(3x+1)?

OpenStudy (anonymous):

It was a question I worked with recently, and couldn't solve..

OpenStudy (mimi_x3):

Just change it to: \[(3x+1)^{-1} \] then differentiate it.

OpenStudy (anonymous):

No, it's In (3x+1)^-1

OpenStudy (mimi_x3):

woops, sorry

OpenStudy (anonymous):

same idea, just imagine the thing inside natural log as one whole chunk: y = ln(3x+1)^-1 y'= [1/(3x+1^)-1] (-(3x + 1)^-2) ( 3) y'= (3x+1)(-1/(3x + 1)^2) (3) do you see how you can reduce the 3x + 1 with the 1/(3x + 1)^2 y' = -1/(3x+1) (3) y' = -3/(3x + 1)

OpenStudy (anonymous):

Oh Ok. Thanks! :)